# On the contraction properties of Sinkhorn semigroups

### v2

@@ 4.1 Strong mixing models
As underlined in [fienberg, mosteller] in the context of finite state space models, Sinkhorn iterations of the form (16) preserve the cross product Radon-Nikodym ratio in (38) and, hence, for any $n\geq 0$ we have
$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}).$ (48)
@@
This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [borwein, chen, deligiannidis, franklin]. Recall that the Hilbert contraction coefficient of a Markov transition $\cal Q$} can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq. (40)), hence the exponential convergence of Sinkhorn iterations w.r.t. the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$. An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
From (39), let us observe that
$\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})\geq\jmath({\cal Q})^{2}.$
Thus, for any given convex function $\Phi$ satisfying (12), we can use (40) and (47) to show that
$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
@@
For any convex function $\Phi$ satisfying (12) we have the contraction estimates
$\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$ (50)
The above inequalities remain valid when we replace $(\nu_{V},\pi_{2n})$ by $(\lambda_{U},\pi_{2n+1})$. If we assume that $\jmath({\cal Q})>0$, then $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ and, in addition, there exists some $c>0$ such that
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~\raisebox{2.0pt}{$\chi$}({\cal S})^{2n}\quad\text{for any $n\geq 0$.}$ (51)
@@
The estimates (50) are a direct consequence of (47) combined with the Markov transport formulae
$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\cal S}_{2(n+1)},\pi_{2n+1}{\cal S}_{2(n+1)}),\quad\text{and}$
$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
which are themselves easily obtained from expressions (16) and (22). The proof of (51) follows exactly the same arguments as the proof of the inequality (71) in Theorem 4.13, hence it is skipped.
@@
Under the assumptions of Theorem 4.1, we have the entropy exponential decay estimates
$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\raisebox{2.0pt}{$\chi$}({\cal S})^{2n},\quad\text{for any $n\geq 0$}.$
@@
Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [dg-01]. For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself. The choice ${\cal Q}(x,dy)~\propto~e^{-\delta\,\|y-b(x)\|}~dx$ yields
@@
$\jmath({\cal Q})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
@@
In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [aronson], Nash [nash] and Varopoulos [varopoulos] on Gaussian estimates for heat kernels on manifolds. Nevertheless, the above condition is not met for Gaussian transitions.
4.2 Gibbs-Loop Sinkhorn semigroups
The $r$-sub-level set of a positive function $\varphi\in{\cal B}(\mathbb{X})$ is denoted as
$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~:~\varphi(x)\leq r\right\}.$ (52)
When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$.
@@ 4.2 Gibbs-Loop Sinkhorn semigroups
Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [dpa] and references therein). In the context of Gibbs-loop semigroups (23), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
Condition 4.3 (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
@@
There exist some parameters $\epsilon\in]0,1[$ some constant $c>0$ and some integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and for any $n\geq n_{0}$ we have the estimates
${\cal S}_{2n}^{\circ}(\psi)\leq\epsilon~\psi+c\quad\mbox{and}\quad{\cal S}_{2n+1}^{\circ}(\varphi)\leq\epsilon~\varphi+c.$ (53)
There also exists some $r_{0}\geq 1$ and some function
[…]
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{2n+1}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}^{\circ}_{2n}\|_{\sf tv}\leq 1-\varepsilon(r).$ (54)
The inequality (54) is satisfied if, and only if, there exists some ${\cal X}\in{\cal M}_{1}(\mathbb{X})$ (respectively, ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$) that may depend on $(n,r,(x_{1},x_{2}))$ (respectively, on $(n,r,(y_{1},y_{2}))$) satisfying, for any $i\in\{1,2\}$, the minorization conditions
${\cal S}^{\circ}_{2n}(y_{i},dy)\geq\varepsilon(r)~{\cal Y}(dy)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}(x_{i},dx)\geq\varepsilon(r)~{\cal X}(dx).$ (55)
The above assertion is rather well known and it can be checked using (33). Note that the drift condition (53) is trivially met for the unit functions $\varphi=1=\psi$ but in this case for any $r\geq 1$ we have $C_{\varphi}(r)=\mathbb{X}$ and $C_{\psi}(r)=\mathbb{Y}$. In that case, the inequalities in (55) are uniform minorization conditions over $\mathbb{X}$ and $\mathbb{Y}$.
@@
Let ${\cal B}_{\infty}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the subalgebra of locally bounded and uniformly positive functions $\varphi$ that grow at infinity; that is, $\sup_{C}\varphi<\infty$ for any compact set $C\subset\mathbb{X}$, and for any
$r\geq\varphi_{\star}:=\inf_{\mathbb{X}}\varphi>0$ (56)
@@
the $r$-sub-level set $C_{\varphi}(r)$ is a non empty compact subset. We denote by ${\cal B}_{0}(\mathbb{X}):=\{1/\varphi~:~\varphi\in{\cal B}_{\infty}(\mathbb{X})\}\subset{\cal B}_{b}(\mathbb{X})$ the sub-algebra of (bounded) positive functions, locally lower bounded and that vanish at infinity. When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ the $r$-sub-level sets $C_{\varphi}(r)$ and $C_{\psi}(r)$ are compact and the inequality (55) yields local minorization conditions.
@@
The assumption $\varphi\wedge\psi\geq 1$ is only made to simplify the presentation. Indeed, for any function $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ satisfying (53), the rescaled functions $\varphi/\varphi_{\star}\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi/\psi_{\star}\in{\cal B}_{\infty}(\mathbb{Y})$ also satisfy (53).
The drift conditions in (53) ensure that for any $n\geq n_{0}$ we have
@@
${\cal S}^{\circ}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\varphi}(\mathbb{X}).$
Whenever Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, it can be easily shown that
$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.\$(57)
These estimates are direct consequences of the drift conditions (53). A detailed proof is provided in Section E on page E.
4.3 Weighted norm contractions
@@
The drift conditions (53) combined with the local minorization conditions (54) provide a way to control the contraction coefficients of Sinkhorn transitions. More precisely, we have the following estimate for the contraction coefficients of the forward-backward Markov transitions ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$.
@@
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied some integer $n_{0}\geq 1$, then there exist parameters $\kappa>0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (53) and on the function $\varepsilon(r)$ in (54) and such that
$\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\cal S}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (58)
Theorem 4.4 is a straightforward consequence of Theorem C.2 in Appendix C.
The contraction estimate (58) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $n_{0}\leq p\leq n$ we readily obtain
@@
with the parameters $\rho\in[0,1[$ and $\kappa>0$ as in (58).
@@
$\|{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{1+n-n_{0}}~\left(\psi_{\kappa}(y_{1})+\psi_{\kappa}(y_{2})\right).$
@@
The proximal sampler discussed in Remark 2.1 has a Markov transition ${\cal S}^{\circ}_{1}$ with invariant measure $\lambda_{U}$. The weighted-norm contraction properties of these models are simply based on the existence of a Lyapunov function $\varphi$ satisfying the r.h.s. condition in (53) with $n=0$. As in (54) we also require a local minorization condition $\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{1}\|_{\sf tv}\leq 1-\varepsilon(r)$ on the sub-level sets $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$ of the Lyapunov function to ensure that $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{1})<1$, for some parameter $\kappa\geq 0$ (see for instance Theorem C.2).
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may not be easy to test – except in the case of Gaussian and linear models. Several Lyapunov functions in this context are designed in Section 5.8. A stronger (but easier to test) condition can be provided in terms of the transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
Condition 4.7 (${\cal D}_{\varphi,\psi}({\cal S})$).
There exists some parameter $\epsilon\in]0,1[$, some constant $c>0$ and some integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$ we have the estimates
${\cal S}_{2n}(\psi)\leq\epsilon~\varphi+c\quad\mbox{and}\quad{\cal S}_{2n+1}(\varphi)\leq\epsilon~\psi+c.$ (60)
There also exists some $r_{0}\geq 1$ and some function
[…]
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (61)
When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, for any $n\geq n_{0}$ we readily see that
${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y}).$
@@
$\pi_{2n+1}(\varphi)=\nu_{V}{\cal S}_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c$ (62)

### v3

@@ 4.1. Strong mixing models
As underlined in [fienberg, mosteller] in the context of finite state space models, Sinkhorn iterations of the form (2.2) preserve the cross product Radon-Nikodym ratio in (3.8) and, hence, for any $n\geq 0$ we have
(4.1) $\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{S}}_{0})=\hbar({\mathcal{Q}}).$
@@
This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [borwein, chen, deligiannidis, franklin]. Recall that the Hilbert contraction coefficient of a Markov transition $\mathcal{Q}$} can be explicitly computed in terms of $\hbar({\mathcal{Q}})$ (see Eq. (3.10)), hence the exponential convergence of Sinkhorn iterations w.r.t. the Hilbert metric takes place if, and only if, $\hbar({\mathcal{Q}})>0$. An elementary proof of this property is presented in Proposition B.1 (in Appendix B). From (3.9), let us observe that
$\jmath({\mathcal{S}}_{n})\wedge\jmath({\mathcal{Q}})\geq\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{Q}})\geq\jmath({\mathcal{Q}})^{2}.$
Thus, for any given convex function $\Phi$ satisfying (1.12), we can use (3.10) and (3.17) to show that
$\hbar({\mathcal{Q}})>0\quad\text{if and only if}\quad\jmath({\mathcal{Q}})>0,$
@@
For any convex function $\Phi$ satisfying (1.12) we have the contraction estimates
(4.3) $\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})~D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2}~D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})~D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2}~D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$
The above inequalities remain valid when we replace $(\nu_{V},\pi_{2n})$ by $(\lambda_{U},\pi_{2n+1})$. If we assume that $\jmath({\mathcal{Q}})>0$, then $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})<1$ and, in addition, there exists some $c>0$ such that
(4.4) $\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2n}\quad\text{for any $n\geq 0$.}$
@@
The estimates (4.3) are a direct consequence of (3.17) combined with the Markov transport formulae
$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\mathcal{S}}_{2(n+1)},\pi_{2n+1}{\mathcal{S}}_{2(n+1)}),\quad\text{and}$
$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\mathcal{S}}_{2n+1},\nu_{V}{\mathcal{S}}_{2n+1}),$
which are themselves easily obtained from expressions (2.2) and (2.8). The proof of (4.4) follows exactly the same arguments as the proof of the inequality (4.24) in Theorem 4.8, hence it is skipped.
@@
If we assume that $\jmath({\mathcal{Q}})>0$, then $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})<1$ and, in addition, there exists some $c>0$ such that
$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})\leq c~\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2n},\quad\text{for any $n\geq 0$}.$
@@
Several examples satisfying condition $\jmath({\mathcal{Q}})>0$ are discussed in Section 2.4 in [dg-01]. For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself. The choice ${\mathcal{Q}}(x,dy)~\propto~e^{-\delta\,\|y-b(x)\|}~dx$ yields
@@
$\jmath({\mathcal{Q}})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
@@
In continuous time settings, condition $\jmath({\mathcal{Q}})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [aronson], Nash [nash] and Varopoulos [varopoulos] on Gaussian estimates for heat kernels on manifolds. Nevertheless, the above condition is not met for Gaussian transitions.
4.2. Gibbs-Loop Sinkhorn semigroups
For a positive function $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $r>0$, we define its $r$-sublevel set by
(4.5) $C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~:~\varphi(x)\leq r\right\}.$
When there is no ambiguity, we write $C_{\varphi}(r)=\{\varphi\leq r\}$.
@@ 4.2. Gibbs-Loop Sinkhorn semigroups
Lyapunov functions combined with local minorization conditions are classical tools in the analysis of time-inhomogeneous Markov chains; see, for instance, [dpa] and the references therein. In the context of Gibbs-loop semigroups (2.9), such conditions are formulated using Lyapunov-type functions $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $\psi\in{\mathcal{B}}(\mathbb{Y})$ satisfying $\varphi\wedge\psi\geq 1$, as described below.
@@
Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$: There exist $\epsilon\in(0,1)$, a constant $c>0$, and an integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and, for all $n\geq n_{0}$,
(4.6) ${\mathcal{S}}_{2n}^{\circ}(\psi)\leq\epsilon,\psi+c,\qquad{\mathcal{S}}_{2n+1}^{\circ}(\varphi)\leq\epsilon,\varphi+c.$
Moreover, there exist $r_{0}\geq 1$ and a function
$\varepsilon:[r_{0},\infty[\rightarrow]0,1]$
such that, for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$, and $n\geq n_{0}$,
(4.7) $\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}^{\circ}_{2n+1}\|{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\mathcal{S}}^{\circ}_{2n}\|{\sf tv}\leq 1-\varepsilon(r).$
Condition (4.7) holds if and only if there exist probability measures ${\mathcal{X}}\in{\mathcal{M}}_{1}(\mathbb{X})$ and ${\mathcal{Y}}\in{\mathcal{M}}_{1}(\mathbb{Y})$, possibly depending on $(n,r,(x_{1},x_{2}))$ and $(n,r,(y_{1},y_{2}))$, respectively, such that, for $i\in{1,2}$,
(4.8) ${\mathcal{S}}^{\circ}_{2n}(y_{i},dy)\geq\varepsilon(r)~{\mathcal{Y}}(dy)\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}^{\circ}(x_{i},dx)\geq\varepsilon(r)~{\mathcal{X}}(dx).$
The drift condition (4.6) is trivially satisfied for the constant functions $\varphi=1=\psi$. In this case, $C_{\varphi}(r)=\mathbb{X}$ and $C_{\psi}(r)=\mathbb{Y}$ for all $r\geq 1$, and (4.8) reduces to a uniform minorization condition over the whole state spaces.
@@
Let ${\mathcal{B}}_{\infty}(\mathbb{X})\subset{\mathcal{B}}(\mathbb{X})$ denote the subalgebra of locally bounded, uniformly positive functions $\varphi$ that grow at infinity, in the sense that $\sup_{C}\varphi<\infty$ for any compact set $C\subset\mathbb{X}$ and that, for any
(4.9) $r\geq\varphi_{\star}:=\inf_{\mathbb{X}}\varphi>0,$
the sublevel set $C_{\varphi}(r)$ is nonempty and compact. We further define
${\mathcal{B}}_{0}(\mathbb{X}):=\{1/\varphi~:~\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})\}\subset{\mathcal{B}}_{b}(\mathbb{X})$
@@
the algebra of bounded, positive functions that are locally lower bounded and vanish at infinity. If $\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})$ and $\psi\in{\mathcal{B}}_{\infty}(\mathbb{Y})$, then the sets $C_{\varphi}(r)$ and $C_{\psi}(r)$ are compact, and (4.8) yields genuine local minorization conditions.
@@
The assumption $\varphi\wedge\psi\geq 1$ is made solely for notational convenience. Indeed, if $\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})$ and $\psi\in{\mathcal{B}}_{\infty}(\mathbb{Y})$ satisfy (4.6), then the rescaled functions $\varphi/\varphi_{\star}$ and $\psi/\psi_{\star}$ also satisfy the same drift inequalities.
The drift conditions (4.6) imply that, for all $n\geq n_{0}$,
@@
${\mathcal{S}}^{\circ}_{2n}\left({\mathcal{B}}_{\psi}(\mathbb{Y})\right)\subset{\mathcal{B}}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}^{\circ}\left({\mathcal{B}}_{\varphi}(\mathbb{X})\right)\subset{\mathcal{B}}_{\varphi}(\mathbb{X}).$
Moreover, whenever ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds,
(4.10) $\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.\$
These bounds follow directly from the drift inequalities (4.6); a detailed proof is given in Section E.
4.3. Weighted norm contractions
@@
Combining the drift conditions (4.6) with the local minorization conditions (4.7) allows us to control the contraction coefficients of the Sinkhorn transitions. More precisely, we obtain the following uniform contraction estimate for the forward–backward Markov kernels ${\mathcal{S}}^{\circ}_{2n}$ and ${\mathcal{S}}^{\circ}_{2n+1}$.
@@
Assume that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds for some $n_{0}\geq 1$. Then there exist parameters $\kappa>0$ and $\rho\in[0,1[$ depending only on $(\epsilon,c)$ in (4.6) and on $\varepsilon(r)$ in (4.7) and such that
(4.11) $\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\mathcal{S}}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\mathcal{S}}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
Theorem 4.3 is an immediate consequence of Theorem C.1 in Appendix C.
The contraction estimate (4.11) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\mathcal{M}}_{1,\psi}(\mathbb{Y})$ and $n_{0}\leq p\leq n$ we readily obtain
@@
with the parameters $\rho\in[0,1[$ and $\kappa>0$ as in (4.11).
@@
$\|{\mathcal{S}}^{\circ}_{2(n_{0}-1),2n}(y_{1},\mbox{\LARGE.})-{\mathcal{S}}^{\circ}_{2(n_{0}-1),2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{1+n-n_{0}}~\left(\psi_{\kappa}(y_{1})+\psi_{\kappa}(y_{2})\right).$
@@
The proximal sampler discussed in Remark 2.1 has a Markov transition ${\mathcal{S}}^{\circ}_{1}$ with invariant measure $\lambda_{U}$. The weighted-norm contraction properties of these models are simply based on the existence of a Lyapunov function $\varphi$ satisfying the r.h.s. condition in (4.6) with $n=0$. As in (4.7) we also require a local minorization condition $\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}^{\circ}_{1}\|_{\sf tv}\leq 1-\varepsilon(r)$ on the sub-level sets $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$ of the Lyapunov function to ensure that $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\mathcal{S}}^{\circ}_{1})<1$, for some parameter $\kappa\geq 0$ (see for instance Theorem C.1).
Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\mathcal{S}}^{\circ}_{2n}$ and ${\mathcal{S}}^{\circ}_{2n+1}$, which may not be easy to test – except in the case of Gaussian and linear models. Several Lyapunov functions in this context are designed in Section 5.8. A stronger (but easier to test) condition can be provided in terms of the transitions ${\mathcal{S}}_{2n}$ and ${\mathcal{S}}_{2n+1}$.
Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ There exists $\epsilon\in]0,1[$, a constant $c>0$ and an integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$,
(4.13) ${\mathcal{S}}_{2n}(\psi)\leq\epsilon~\varphi+c,\qquad{\mathcal{S}}_{2n+1}(\varphi)\leq\epsilon~\psi+c.$
Moreover, there exists some $r_{0}\geq 1$ and a function
$\varepsilon:[r_{0},\infty[\rightarrow]0,1]$
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$,
(4.14) $\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\mathcal{S}}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$
When Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is met, for any $n\geq n_{0}$ we readily see that
${\mathcal{S}}_{2n}\left({\mathcal{B}}_{\psi}(\mathbb{Y})\right)\subset{\mathcal{B}}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}\left({\mathcal{B}}_{\varphi}(\mathbb{X})\right)\subset{\mathcal{B}}_{\psi}(\mathbb{Y}).$
@@
(4.15) $\pi_{2n+1}(\varphi)=\nu_{V}{\mathcal{S}}_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c$